OpenStudy (anonymous):

Consider the function f(x)=4x^2-3x^4. Is the function even, odd, or neither? Evaluate the function at the given values of the independent variable and simplify. f(-3), f(-x)

OpenStudy (jamesj):

A function is odd if f(x) = -f(-x) A function is even if .... Check whether or not the function you've written down satisfies either definition.

OpenStudy (anonymous):

f(-x) = -f(x) means odd f(-x) = f(x) means even the function is even. f(-3) = 4 * (-3)^2 - 3 * (-3)^4 = -207 f(-x) = 4 * (-x)^2 - 3*(-x)^4 -> 4x^2 -3x^4 which is not surprising since the function is even :)